OpenStudy (anonymous):

10w + 3z = 29 when z = 3

OpenStudy (anonymous):

substitute 10w + 3(3) = 29 10w + 9 = 29 subtract 9 10w +9 -9 = 29 - 9 10w = 20 divide by 10 w = 2

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

but why did u divdd by 10

OpenStudy (anonymous):

divide both sides by 10 to isolate w on left-wide 10w/10 = 20/10 w = 2

OpenStudy (anonymous):

o ok